OpenStudy (anonymous):

write an equation of the line that passes through the given points. (-1,3), (2,9) 9-3 ---- 2+1 = 6/3=2 y=2x is this correct?

OpenStudy (maheshmeghwal9):

no u r not right easy to check if u put -1 in ur line's equation u must get 3 but u r getting -2 so add 5 to ur equation.

OpenStudy (maheshmeghwal9):

actually u have missed the intercept part:/

OpenStudy (maheshmeghwal9):

y=mx+c y=2x+5.

OpenStudy (maheshmeghwal9):

now this is correct:)

OpenStudy (anonymous):

could you arrange it so i could understand better..

OpenStudy (anonymous):

y - y1 = m ( x - x1)

OpenStudy (maheshmeghwal9):

y = 2x is wrong u must add intercept here if u calculate easily in my first comment as i said above u would get intercept as 5 so ur equation will become y= 2x+5 now this is right:) gt it ?

OpenStudy (maheshmeghwal9):

http://www.wolframalpha.com/input/?i=y%3D+2x%2B5